# Factual Knowledge in Language Models: Robustness and Anomalies under Simple Temporal Context Variations

### v3

@@ 4.2 Robustness and Anomalies
Figure 6 reports this transfer proportion from granularity $B$ to $A$ for the "plain text" format (top) and "instruction" format (bottom). On average, LMs failed to generalize their knowledge to other granularities in 38% of cases (average of all non-diagonal cells), which is surprisingly high given their perfect score on the starting granularity $Y$. Details for each model are available in Appendix C.1. Performance varies across LMs. For example, for the most robust model, gemma-2-27b-it, the transition from $B=\mbox{Y}$ to $A=\mbox{YM}$ is successful in 70±6% of cases, and the win rates for other transitions range between 63±7% and 80±6%. Consistent with previous observations, the LM with the highest overall success rate (Llama-3.1-70B-Instruct) performs worse in terms of transfer between granularities, with proportions ranging from 39±8% to 65±10%. No LM achieves perfect transitions for any pair of granularities. The "instruction" format (Figure 6, bottom) improves transfer, with an average proportion now at 68%. The improvements are primarily driven by transfers from a coarser granularity to a finer granularity (lower diagonal of the matrix).
@@
For exploratory purposes, we investigated whether including explanations about temporal concepts in the LMs’ prompts could help them better transfer knowledge from one temporal granularity to another. To evaluate this, two prompts were prefixed to each TimeStress statement. The first explains the hierarchical nature of dates (i.e., a year consists of months, and a month consists of days), while the second is more direct and explains how knowledge of a temporal fact can be generalized from one granularity to another. Details of these prompts are provided in Appendix C.2.
We recalculated the transfer proportions between granularities using the same 5 LMs as in Figure 6. The two explanatory prompts improved generalization in the "plain text" format from 62% to 68%. However, no substantial gain compared to not using an explanatory prompt was observed when using the "instruction" format.
5 Discussion and Conclusion
This study examined the robustness of language models (LMs) to simple temporal variations in factual knowledge. It assessed their ability to distinguish correct from incorrect temporal contexts based on two factors: the distance of contexts from the validity period of facts and their granularity. To facilitate this, the TimeStress dataset was introduced, featuring high-quality statements on popular temporal facts from Wikidata, and enabling the evaluation of 18 LMs of varying sizes and families.
@@ 5 Discussion and Conclusion
The results revealed that the best-performing LM was robust for only 6% of the studied facts, exhibiting critical errors uncommon to humans, which we frame as anomalies. These errors consist in a susceptibility to easy incorrect contexts and a limited knowledge generalization across granularities. Notably, these findings held true regardless of whether the LM was pretrained or instruction-tuned, and whether the statements were presented in an instruction or plain format. This highlights the limits of current LMs in temporal representation.
It is worth noting that since the studied temporal facts are popular, these results likely represent an upper bound of LMs’ performance on the general population of facts, given the strong link between knowledge popularity and its likelihood of being learned by LMs Kandpal et al. (2023); Kang and Choi (2023).
@@ Appendix A TimeStress: Details of the Construction Process
First, the dataset focuses on past facts valid strictly before 2021, ensuring that they are historical (not valid at the present) events for all recent LMs. TimeStress includes high-quality statements that are consistent with the facts and exhibit linguistic diversity to avoid biases stemming from a limited variety of questions. The statements are carefully selected to minimize typographical errors, verbs are systematically conjugated in the past tense, and future dates beyond 2020 are excluded to avoid absurd questions such as "In 2052, who was the president of the USA?". The dataset covers a diverse set of 86 relations to reduce biases associated with a restricted range. The targeted facts are popular, which is crucial for the objective of our study and essential for evaluating the generalization of knowledge across different granularities—a task that becomes challenging if the LMs are unfamiliar with the facts. All facts are valid over a single validity period, ensuring that all contexts outside the validity period can be considered incorrect. Additionally, to ensure fairness, each granularity (Y, YM, YMD) has an equal number of correct and incorrect temporal contexts for all facts. Finally, the number of correct and incorrect contexts is sufficiently large to make it nearly impossible for a random model to robustly know a fact by chance.
The creation process for the TimeStress dataset was carefully designed to meet the properties described above, thereby effectively supporting the claims of this paper. This process consists of three main steps. First, an initial collection of 2,098 popular temporal facts is performed from Wikidata for inclusion in TimeStress. Second, questions are generated from these quintuplets using GPT-4o, accompanied by a quality evaluation to ensure high-quality questions. Finally, for each fact, correct and incorrect temporal contexts are identified and integrated into the questions to produce statements.
@@ A.1 Quintuplet Collection Process
This source also provides a measure of an entity’s popularity, defined as the median number of human visits to the Wikipedia article associated with that entity during the year 2020. This measure is used to define the popularity of a quintuplet, calculated as the geometric mean of the popularity of its object and subject. Figure 12 demonstrates the effectiveness of this popularity measure in identifying facts known by LMs, illustrating that the likelihood of LMs knowing a fact increases with its popularity.
@@
Only quintuplets with a popularity measure of at least 90,000333This threshold was determined by gradually lowering the threshold from 150,000 in steps of 10,000 until the number of retrieved facts exceeded 2,000. and a validity period longer than three years are retained.
The final result is a dataset comprising the 2,098 most popular facts from Wikidata, with 1,910 unique entities, 1,435 unique subjects, 1,151 unique objects, and 86 relations, forming a well-diversified set of temporal facts.
@@ A.2 Quintuplet Verbalization
The process of verbalizing quintuplets into natural language questions is carried out using GPT-4o. The prompt, adapted from Ammar Khodja et al. (2024) (Appendix B), was modified to generate questions instead of declarative sentences. The adapted system prompt instructs GPT-4o to take a tuple (subject, relation, object, timestamp) and generate four linguistically diverse questions. For example, for the input (British India, capital, Kolkata, 1929), a possible question could be: "In 1929, what was the capital of British India? Kolkata". The questions must adhere to specific guidelines: they must be in the past tense, begin with the year followed by a comma, and end with the answer. The questions should focus on the object, be simple and concise, and avoid any details that could simplify the answer.
@@ A.4 Test Generation
For each context of granularity Y, a context of granularity YM is chosen by randomly selecting a month within the year. Similarly, for each context of granularity YM, a context of granularity YMD is chosen by randomly selecting a day within the previously selected YM context. This creates a hierarchical relationship between the different granularities (e.g., 2020, 2020-03, 2020-03-24), enabling reasonable comparisons in terms of win rates and robustness, as they share the same year and/or month. All contexts are now classified as correct, incorrect, or transitional (cf. Section 3.1).
@@ Appendix B Vulnerability to Easy Incorrect Contexts: Analysis of Results at Different Win Rate Thresholds
In Section 4.2, we demonstrated that LMs, even when they almost perfectly know a fact, often fail to achieve "perfect knowledge" due to their vulnerability to easy contexts that are far outside the validity period (Table 4). We defined a fact as being almost perfectly known by an LM if its win rate exceeded 95% but did not reach 100%. In this section, we extend this analysis by experimenting with different win rate thresholds to observe how the distribution of incorrect contexts favored over correct contexts evolves as the threshold approaches 1.
@@
Figure 7: Proportion of incorrect contexts favored over correct contexts that are beyond a relative distance $\alpha$ from the validity period, when the win rate exceeds the threshold, for the 5 most robust LMs. Experiments were conducted with granularity Y.
@@ Appendix C Generalization of Knowledge Across Granularities
C.1 Generalization Matrices for Each LM
In Section 4.2, we explored the ability of language models to generalize their temporal knowledge from one granularity to another. We provided two matrices (one for instruction-based questions and one for plain text questions) containing the generalization rate between each granularity pair averaged over the 5 most robust LMs. Complementing these average performances, the generalization rate matrices for individual models are presented in Figure 8.
@@ C.1 Generalization Matrices for Each LM
Figure 8: Generalization matrics between pairs of granularities on the 5 most robust LMs. In the first row, the statements are presented in a plain format, and in the second row, they are presented in a instruction format.
C.2 Explanatory Prompts
@@
In addition to the average performance in the 4.2 section, figure 9 shows the average generalization matrices across the same 5 models as in figure 6, using plain text and an instruction format.
(a) Prompt 1, plain text
@@
(e) No explanation prompt, plain text
@@ Appendix D Conditional Probability Calculations in LMs
Given that different tokenizers split a text differently, we require a universal algorithm to calculate at best the probability of generating a text given a prompt, even when the end of the prompt might be in the middle of a token.
@@
    Tokenize $A+B$ into a sequence of tokens $s=(t_{1},t_{2},\dots,t_{n})$555$+$ represents the string concatenation operation..
@@ Appendix E Supplementary Results
  • Figure 14 displays the relations that were most robustly known on average by the studied LMs (plain text format statements).
  • The pretrained models and their respective instruction-tuned versions are compared using matches between their win rate scores $\mathcal{W}$ on each fact (Figure 12). The instruction-tuned models are more informed than the pretrained models on questions using win rate as a metric. Regarding robustness, the vast majority of matches resulted in ties.
Figure 11: Relationship between fact popularity and robustness metric calculated across granularities Y, YM, YMD.
Figure 12: Comparison of the performance of pretrained LMs and their respective instruction-tuned version by comparing their win rate on each fact in TimeStress.

### v6

@@ 4.2 Robustness and Anomalies
Figure 6 reports this transfer proportion from granularity $B$ to $A$ for the "raw text" format (left) and "instruction" format (right). On average, for the "raw text" format, LMs failed to generalize their knowledge to other granularities in 28% of cases (1 - average of all non-diagonal cells), which is surprisingly high given their perfect score on the starting granularity $Y$. Details for each model are available in Appendix C.2. Performance varies across LMs. For example, for the most robust model, gemma-2-27b-it, the transition from $B=\mbox{Y}$ to $A=\mbox{YM}$ is successful in 74±5% of cases, and the win rates for other transitions range between 68±6% and 88±5%. The general trend is that LMs fail more in transitions from coarse to fine granularities. No LM achieves perfect transitions for any pair of granularities. There are slight variations between the instruction (Figure 6, right) and raw formats, but the average success rate is nearly identical.
The possibility that poor knowledge propagation between granularities could be due to LMs’ ignorance of the validity period boundaries [^6] In this case, robustness was achieved only by chance.. This was confirmed in a similar analysis that takes context position into account (Appendix C.1). Indeed, consistency between granularities approaches perfect consistency as the context moves away from the validity period. However, perfect consistency is never reached; which reminds us of the vulnerability of LMs to easy incorrect contexts.
@@
For exploratory purposes, we investigated whether including explanations about temporal concepts in the LMs’ prompts could help them better transfer knowledge from one temporal granularity to another. To evaluate this, two prompts were prefixed to each TimeStress statement. The first explains the hierarchical nature of dates (i.e., a year consists of months, and a month consists of days), while the second is more direct and explains how knowledge of a temporal fact can be generalized from one granularity to another. Details of these prompts are provided in Appendix C.3. We recalculated the transfer proportions between granularities using the same 5 LMs as in Figure 6. The two explanatory prompts improved generalization in the "raw text" format from 73% to 76%. However, no substantial gain compared to not using an explanatory prompt was observed when using the "instruction" format.
Other observations.
There is a positive correlation between the popularity of a fact and the robustness and win rate of LMs on it. Interestingly, LMs are robust on globally different facts. Indeed, a pair of LMs shares, on average, 11% of facts on which they are robust. This proportion reaches 31% when limited to the 5 most robust LMs. However, only 34 facts out of 384 (8.9%) are robust at the same time in these LMs. Furthermore, the longer a fact’s validity period, the higher the win rate (on the 5 most robust LMs). This statistically significant correlation [^7] The null hypothesis is the absence of correlation. is intriguing because it appears that the difficulty of situating a fact in time is the same whether it has a duration of 3 years or 30 years. One possible explanation is that facts with longer validity periods are more stable and unique (i.e., there are no alternative objects "o" for the same subject-relation pair "s,r"), so LMs can learn them without confusion or contradiction. However, this explanation is contradicted by another observation: when there are more alternative objects "o" for a given (s,r) pair, the win rate and robustness actually increase, not decrease. This contradiction raises the question of how to explain the observed phenomenon. Finally, the further a fact’s validity period is from the present, the less robust the LMs are on it, with lower win rates as well. More details are in Appendix E.
5 Experimental Protocol: Motivations
There are seemingly more "natural" approaches for probing factual knowledge in language models, such as the evaluation protocols used in LAMA Petroni et al. (2019), TriviaQA Joshi et al. (2017), KAMEL Kalo and Fichtel (2022), and BEAR Wiland et al. (2024). Instead of comparing probabilities across several temporal contexts, one could ask the LM to answer temporally contextual questions such as “In 2011, who was the president of the US?”, and evaluate the LM based on the generated answers. However, our experimental protocol was preferred for several reasons.
First, our setup–where the LM must distinguish between statements with correct and incorrect temporal contexts by assigning probabilities–allows to target specific facts without ambiguity, even in the case of non-functional relations, such as "shares a border with," where a subject-relation pair can have multiple valid objects. In generation-based settings, an LM may produce one or several correct answers, or even off-topic outputs, making evaluation less reliable and direct comparison across LMs more difficult. This is especially true given that classical generation-based metrics, such as ROUGE Lin (2004), can underestimate performance. Sometimes, the set of all correct answers is difficult to enumerate due to the vagueness of the relation (e.g., does asking for the borders of a country include continents and oceans?) and due to the sometimes large number of ways of expressing an answer.
Additionally, our evaluation protocol is efficient and scalable, as it does not require generation or answer validation.
Given the imperfections of other evaluation protocols, it would have been difficult to defend our claims–especially those involving sensitive metrics like robustness and the study of rare LM errors–if our results could be attributed to limitations of the evaluation method itself.
6 Conclusion
@@
This study examined the robustness of LMs to simple temporal variations in factual knowledge. It assessed their ability to distinguish correct from incorrect temporal contexts based on two factors: the distance of contexts from the validity period of facts and their granularity. To facilitate this, the TimeStress dataset was introduced, featuring high-quality statements on popular temporal facts from Wikidata (according to a popularity index) and enabling the evaluation of 18 LMs of varying sizes and families. The results revealed that the best-performing LM was robust for only 11% of the studied facts, exhibiting errors, certainly rare, but critical that are uncommon to humans, which we frame as anomalies. These errors consist of a susceptibility to easy incorrect contexts and imperfect knowledge generalization across granularities. Notably, these findings held true regardless of whether the LM was pretrained or instruction-tuned, and whether the statements were presented in an instruction or raw format. This highlights the limits of current LMs in temporal representation. It is worth noting that since the studied temporal facts are relatively popular, these results likely represent an upper bound of LMs’ performance on the general population of facts, given the strong link between knowledge popularity and its likelihood of being learned by LMs Kandpal et al. (2023); Kang and Choi (2023).
@@
Second, the results of our study are limited to the format of the statements we chose, i.e., a temporal context followed by a question and an answer. It is possible that LMs would perform better in a different format. However, their current limitations on our data are already problematic.
@@ Appendix A TimeStress: Details of the Construction Process
First, the dataset focuses on past facts valid strictly before 2021, ensuring that they are historical (not valid at the present) events for all recent LMs. TimeStress includes high-quality statements that are consistent with the facts and exhibit linguistic diversity to avoid biases stemming from a limited variety of questions. The statements are carefully selected to minimize typographical errors, verbs are systematically conjugated in the past tense, and future dates beyond 2020 are excluded to avoid absurd questions such as "In 2052, who was the president of the USA?". The dataset covers a diverse set of 86 relations to reduce biases associated with a restricted range. The targeted facts are popular, essential for evaluating the generalization of knowledge across different granularities—a task that becomes challenging if the LMs are unfamiliar with the facts. All facts are valid over a single validity period, ensuring that all contexts outside the validity period can be considered incorrect. Additionally, to ensure fairness, each granularity (Y, YM, YMD) has an equal number of correct and incorrect temporal contexts for all facts. Finally, the number of correct and incorrect contexts is sufficiently large to make it nearly impossible for a random model to be robust on any fact by chance.
The creation process for the TimeStress dataset was carefully designed to meet the properties described above, thereby effectively supporting the claims of this paper. This process consists of three main steps. First, an initial collection of 2,098 temporal facts is performed from Wikidata for inclusion in TimeStress. Second, questions are generated from these quintuplets using GPT-4o, accompanied by a quality evaluation to ensure high-quality questions. Finally, for each fact, correct and incorrect temporal contexts are identified and integrated into the questions to produce statements.
@@ A.1 Quintuplet Collection Process
This source also provides a measure of an entity’s popularity, defined as the median number of human visits to the Wikipedia article associated with that entity during the year 2020. This measure is used to define the popularity of a quintuplet, calculated as the geometric mean of the popularity of its object and subject. Figure 14 demonstrates the effectiveness of this popularity measure in identifying facts on which LMs are robust, illustrating that the likelihood of the robustness of LMs on a fact increases with its popularity.
@@
Only quintuplets with a popularity measure of at least 90,000888This threshold was determined by gradually lowering the threshold from 150,000 in steps of 10,000 until the number of retrieved facts exceeded 2,000. and a validity period strictly longer than three years are retained.
The final result is a dataset comprising the 2,098 most popular facts from Wikidata (according to the popularity index), with 1,910 unique entities, 1,435 unique subjects, 1,151 unique objects, and 86 relations, forming a well-diversified set of temporal facts.
@@ A.2 Quintuplet Verbalization
The process of verbalizing quintuplets into natural language questions is carried out using GPT-4o. The prompt, adapted from Ammar Khodja et al. (2024) (Appendix B), was modified to generate questions instead of declarative sentences. The adapted system prompt instructs GPT-4o to take a tuple (subject, relation, object, timestamp) and generate four linguistically diverse questions. For example, for the input (British India, capital, Kolkata, 1929), a possible question could be: "In 1929, what was the capital of British India? Kolkata". The questions must adhere to specific guidelines: they must be in the past tense, begin with the year followed by a comma, and end with the answer. The questions should focus on the object, be simple and concise, and avoid any detail that could simplify the answer.
@@ A.4 Test Generation
For each context of granularity Y, a context of granularity YM is chosen by randomly selecting a month within the year. Similarly, for each context of granularity YM, a context of granularity YMD is chosen by randomly selecting a day within the previously selected YM context [^10] This sampling does not produce erroneous dates such as February 29 for non-leap years, or April 31.. This creates a hierarchical relationship between the different granularities (e.g., 2020, 2020-03, 2020-03-24), enabling reasonable comparisons in terms of win rates and robustness, as they share the same year and/or month. All contexts are now classified as correct, incorrect, or transitional (cf. Section 3.1).
@@ Appendix B Vulnerability to Easy Incorrect Contexts: Analysis of Results at Different Win Rate Thresholds
In Section 4.2, we demonstrated that LMs, even when they are almost robust on a fact (i.e., a high win rate but inferior to 100%), often fail to achieve robustness due to their vulnerability to easy contexts that are far outside the validity period (Table 4). In this section, we extend this analysis by experimenting with different win rate thresholds to observe how the distribution of incorrect contexts favored over correct contexts evolves as the threshold approaches 100%.
@@
Figure 7: Proportion of incorrect contexts favored over correct contexts that are beyond a relative distance $\alpha$ from the validity period, when the win rate exceeds the threshold, for the 5 most robust LMs. Experiments were conducted with granularity Y. 95% confidence intervals were calculated using bootstrapping.
@@ Appendix C Generalization of Knowledge Across Granularities
C.1 Consistency Across Granularities Based on Relative Distance
In this section, we examine the consistency of LM predictions across different granularities (Y, YM, YMD) as the distance between the tested context and the validity period increases.
To evaluate this, and solely for this section, we introduce a metric called local robustness. Local robustness for a fact, a LM, and a given incorrect context is equal to 1 if all correct contexts are preferred over this incorrect context, and 0 otherwise.
We group all statements in TimeStress according to the relative distance $\alpha$ from their temporal context, and restricting ourselves to the 5 most robust MLs and to the "known" facts [^11] We recall that ”known” in the context of this article means that the ML in question has a robustness equal to 1 on the fact in question, i.e., all correct contexts are preferred to incorrect contexts by the ML. at least on one granularity by these LMs. These statements are categorized according to the interval of which their relative distance $\alpha$ is part. The chosen intervals are $]s,s+\frac{1}{2}]$, where $s$ can take values from $\{-5,-4.5,\dots,4.5\}$. For each interval, the contexts are aligned by fact and by granularity hierarchically (e.g., 2020, 2020-04, 2020-04-23), which is guaranteed to be possible due to the properties of TimeStress (cf. Section 3.2). Local robustness is then calculated for each incorrect context, and the accuracy [^12] Accuracy measures the proportion of identical elements between two vectors, that is, the number of positions where the values are equal, divided by the total number of elements compared. between these measures is computed for all granularity pairs (i.e., Y-YM, Y-YMD, and YM-YMD). These coefficients are averaged across all granularity pairs, all facts, and the 5 most robust LMs, with the results presented in Figure 8.
The results indicate that the inconsistency between granularities is mainly caused by incorrect contexts located at the boundaries of the validity period. As the context moves away from the validity period, the consistency approaches a perfect score of 1 but never reaches it regardless of the ML, the statement type and the $\alpha$ interval used.
Figure 8: For each $\alpha$ segment, the average local robustness correlation across all granularity pairs is calculated over all facts and the 5 most robust LMs.
C.2 Generalization Matrices for Each LM
In Section 4.2, we explored the ability of language models to generalize their temporal knowledge from one granularity to another. We provided two matrices (one for instruction-based questions and one for raw text questions) containing the generalization rate between each granularity pair averaged over the 5 most robust LMs. Complementing these average performances, the generalization rate matrices for individual models are presented in Figure 9.
@@ C.2 Generalization Matrices for Each LM
Figure 9: Generalization matrics between pairs of granularities on the 5 most robust LMs. In the first row, the statements are presented in a raw format, and in the second row, they are presented in a instruction format.
C.3 Explanatory Prompts
@@
In addition to the average performance in the 4.2 section, figure 10 shows the average generalization matrices across the same 5 models as in figure 6, using raw text and an instruction format.
(a) Prompt 1, raw text
@@
(e) No explanation prompt, raw text
@@ Appendix D Conditional Probability Calculations in LMs
Given that different tokenizers split a text differently, we require a universal algorithm to best calculate the probability of generating a text given a prompt, even when the end of the prompt might be in the middle of a token.
@@
    Tokenize $A+B$ into a sequence of tokens $s=(t_{1},t_{2},\dots,t_{n})$131313$+$ represents the string concatenation operation..
@@
(a) Average win rate
(b) Average robustness
Figure 12: Average metrics across all facts in TimeStress for the 18 studied LMs with 95% confidence intervals (determined using bootstrapping).
@@ Appendix E Supplementary Results
  • The average robustness score and win rate across the 18 studied LMs are presented in Figure 12.
@@
  • Figure 16 displays the relations that were most robustly known on average by the studied LMs ("raw text" format statements).
  • Figure 17 shows examples where LMs were vulnerable to easy incorrect contexts.
  • Figure 14 shows the year distribution of temporal contexts across the entire TimeStress dataset.
  • Figure 18 shows the influence of fact distance from the present (here, the year 2021), as well as their durations, on the robustness and win rate of the 5 most robust MLs. The time unit used for both metrics is the year.
Figure 13: Relationship between fact popularity and robustness metric calculated across granularities Y, YM, YMD. The Pearson coefficient is equal $+0.065$ (p-value $<10^{-51}$).
Figure 14: Distribution of the the years of all the temporal contexts in TimeStress.
@@
Figure 17: Examples of vulnerability to easy incorrect contexts for different LMs. The color blue represents the boundaries of the validity period, the color green represents incorrect contexts that are never preferred to correct contexts, and the color red, on the contrary, represents incorrect contexts that were preferred to one or more correct contexts.
(a) Logarithm of the distance of the fact with respect to the present.
(b) Logarithm of the duration of the fact.
Figure 18: The influence of two factors on the robustness and win rate of the 5 most robust LMs. All correlations are statistically significant where the null hypothesis is the absence of linear correlation. Robustness is missing from Figure b because its analysis is not relevant as the duration of a fact is confounded with another variable: the number of matches of a fact. Indeed, the longer a fact is, the more matches it has, and the lower is the robustness.